OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL I think it's a and c but I'm not sure check my answers please How does the graph of f(x)=3(4)^(x-5)+2/3 relate to its parent function? Choose all that apply A the parent function has been translated up. B the parent function has been stretched C the parent function has been translated to the right D the parent function has been compressed.

OpenStudy (anonymous):

12^(x+5)+2/3 is a quardratic function it has been moved down 2/3 units and is strecthed

OpenStudy (anonymous):

So a and c are correct

OpenStudy (anonymous):

it was also translated to the left so its b no c

OpenStudy (anonymous):

if it's x-5 that means its translated right 5 units if it x+5 its translated left 5 units

OpenStudy (anonymous):

a and b

OpenStudy (anonymous):

Now I get a b and c are you sure it's not c @FLAMER123

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

c would mean it's translated to the right but when adding 5 its translated to the left.

OpenStudy (anonymous):

It's subtracting 5 though

OpenStudy (anonymous):

The numbers in this function do the opposite of what they look like they should do. For example, if you have the equation g(x) = (x – 3)2, the graph of f(x)=x2 gets moved to the right three units; in h(x) = (x + 2)2, the graph of f(x)=x2 gets moved to the left two units.

OpenStudy (anonymous):

So it's moving right because it is subtracting

OpenStudy (anonymous):

and its x+2 not x2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So c is correct

OpenStudy (anonymous):

LOL i didn't read your question right LOL sorry yeah your right.

OpenStudy (anonymous):

So the answer is a b and c

OpenStudy (anonymous):

@FLAMER123